### v1

# Instructions for Authors: Title in Title Case

Abstract
Traditional imitation learning focuses on modeling the behavioral mechanisms of experts, which requires a large amount of interaction history generated by some fixed expert. However, in many streaming applications, such as streaming recommender systems, online decision-makers typically engage in online learning during the decision-making process, meaning that the interaction history generated by online decision-makers includes their behavioral evolution from novice expert to experienced expert. This poses a new challenge for existing imitation learning approaches that can only utilize data from experienced experts. To address this issue, this paper proposes an inverse batched contextual bandit (IBCB) framework that can efficiently perform estimations of environment reward parameters and learned policy based on the expert’s behavioral evolution history. Specifically, IBCB formulates the inverse problem into a simple quadratic programming problem by utilizing the behavioral evolution history of the batched contextual bandit with inaccessible rewards. We demonstrate that IBCB is a unified framework for both deterministic and randomized bandit policies. The experimental results indicate that IBCB outperforms several existing imitation learning algorithms on synthetic and real-world data and significantly reduces running time. Additionally, empirical analyses reveal that IBCB exhibits better out-of-distribution generalization and is highly effective in learning the bandit policy from the interaction history of novice experts.
1 Introduction
Traditional imitation learning (IL) focuses on learning the decision-making policy of an experienced and fixed expert by using its historical behaviors. IL utilizes experts’ demonstrations for policy fitting and aims to extract knowledge from experts’ demonstrations to replicate their behaviors or environments. It can be divided into two main categories: Behavior Cloning (BC) [Bain and Sammut, 1995] and Inverse Reinforcement Learning (IRL) [Russell, 1998]. BC approaches aim to learn a policy that directly maps states to actions, while IRL approaches focus on recovering the reward parameters from demonstrations to recover the experts’ policy [Zheng et al., 2022].
A major challenge in imitation learning (IL) is the substantial requirement of expert demonstration data [Abbeel and Ng, 2004, Hussein et al., 2017]. There is a common assumption that the experts are always experienced (see Figure 1). However, in real-world streaming scenarios, the behavior policy of the expert constantly evolves over time, transitioning from a novice expert to an experienced one. Throughout this process, we accumulate a significant amount of expert’s behavioral evolution history. Taking streaming recommender systems as an example [Zhang et al., 2021, 2022], the agent in streaming recommendation needs to make a trade-off between exploitation and exploration during the recommendation process, and continuously and incrementally update its recommendation policy [Chandramouli et al., 2011, Chang et al., 2017, Jakomin et al., 2020].
Figure 1: Imitation Learning (IL) approaches often hold the view that interactions made by experts are experienced, which may not conform to streaming scenarios since even the expert has to evolve from novice to experienced.
In the behavioral evolution history, there exist a significant amount of contradictory data where an expert may take different actions when facing the same context at different time periods. This contradicts the assumption made by traditional imitation learning (IL) approaches regarding the consistency of expert behavior data. Therefore, directly applying existing IL methods to recover the behavior policy from the expert’s behavioral evolution history is not reasonable. For example, behavior cloning (BC) approaches are highly sensitive to distribution drift and contradictory data, and inconsistencies in the expert’s evolution history can lead to a sharp decline in the final recovery performance. Furthermore, some IL approaches rely on large-scale iterations of Metropolis-Hastings sampling to achieve accurate estimation in universal scenarios [Ramachandran and Amir, 2007, Choi and Kim, 2011]. However, these approaches overlook the exploitation-exploration mechanism in expert interactions and demonstrate low sample efficiency when dealing with large-scale interaction data. While recent work by Hüyük et al. [2022] focused on contextual bandits, it cannot be adapted to the batched bandit setting as it assumes the expert updates the policy in a fully-online manner. Additionally, existing inverse bandit approaches typically employ sampling techniques to estimate the likelihood, resulting in low training efficiency. There is an urgent need to develop inverse bandit methods specifically designed for the more general batched bandit setting and to devise novel techniques for recovering parameters tailored for bandit polices.
In this paper, we focus on the more general batched contextual bandit (BCB) setting, and propose inverse batched contextual bandit (IBCB) to overcome several issues mentioned above. To learn from the evolving behaviors made by online decision-making experts with BCB policies, we design IBCB with a unified framework for both deterministic [Han et al., 2020] and randomized [Dimakopoulou et al., 2019] bandit policies. Under this unified framework, we take linear constraints with expectation relaxation in BCB’s exploitation-exploration pairs and batched updating strategy into consideration. We formulate the inverse problem of IBCB into a simple quadratic programming problem without notifying expert’s behavior rewards. We take the assumption of BCB setting, rather than simply cloning behaviors with no assumption or blindly using large-scale samplings. Combining all these improvements together accelerates the train speed of IBCB and insures the robustness of IBCB for out-of-distribution and contradictory data. Due to that original BCB learns the reward parameter through online learning, IBCB can capture both expert’s policy parameters and reward parameters at the same time.
We summarize our major contributions as follows: (1) We define a new inverse bandit problem with behavioral evolution history in the BCB setting; (2) We introduce a unified framework called IBCB, designed for both deterministic and randomized BCB policies, which can efficiently learn from the interaction history data of novice experts without requiring experts’ feedbacks; (3) IBCB is capable of simultaneously learning the parameters of the expert policy and the reward feedback, which outperforms various existing baseline IL approaches in scenarios such as out-of-distribution data and contradictory data, while also demonstrating significantly improved training speed.
2 Related Works
Contextual Bandits have been extensively utilized for solving sequential decision-making problems in online learning [Li et al., 2010, Lan and Baraniuk, 2016, Yang et al., 2021]. Recently, there has been increased research focus on a more general setting called batched contextual bandit (BCB) [Han et al., 2020, Ren and Zhou, 2020, Gu et al., 2021], where actions within the same batch share fixed policy parameters. In this regard, Han et al. [2020] proposed a UCB-like update strategy for BCB, which is applicable to both random and adversarial contextual data.
Imitation Learning (IL) is typically used to learn the expert’s action selection policy or estimate the reward parameter from the expert’s behavior [Abbeel and Ng, 2004, Hussein et al., 2017, Arora and Doshi, 2021, Zheng et al., 2022]. IL can be divided into two categories. The first category is Behavior Cloning (BC) [Bain and Sammut, 1995], which directly constructs a direct mapping from state to action [Abbeel et al., 2010, Osa et al., 2017, Liu et al., 2020]. The second category is the inverse reinforcement learning (IRL) method [Russell, 1998], which attempts to recover environment’s reward parameter from the expert’s behavior, that is, imitating the reward parameters. Existing works mainly study IRL based on Bayesian [Ramachandran and Amir, 2007, Choi and Kim, 2011, Balakrishna et al., 2020], maximum entropy (max-ent) [Ziebart et al., 2008, Fu et al., 2018, Qureshi et al., 2019], and generative [Ho and Ermon, 2016, Chen et al., 2021] approaches. Traditional IL approaches mostly assume that the expert’s behavior is optimal at every step [Ziebart et al., 2008, Hussein et al., 2017, Brown and Niekum, 2019], but ignore that early interactions are likely to be the evolutionary data of novice experts learned to become experienced experts.
Offline Reinforcement Learning (Offline RL) has been widely used to improve the model’s ability to generalize to out-of-distribution data, i.e., extrapolation error [Prudencio et al., 2023, Levine et al., 2020]. There are two main research focuses in existing offline RL. One is to restrict the policy during training to avoid out-of-distribution generalization issues, which can be further subdivided into explicitly restricted policies [Mnih et al., 2015, Fujimoto et al., 2019, Kumar et al., 2020, Kostrikov et al., 2021]and implicitly restricted policies [Peng et al., 2019, Chen et al., 2020]. The other is to allow the learning model to adaptively estimate the uncertainty of the environment or data [Kidambi et al., 2020, Yu et al., 2021]. Unlike Offline RL, which requires knowledge of the reward feedback, the proposed IBCB can learn about reward and expert policy’s parameters from the history of expert behavioral evolution without requiring reward feedback information from the environment.
Figure 2: Inverse bandit problem with behavioral evolution history in BCB setting at step $b$ in $n$-th episode, where the rewards are invisible and our goal is to infer and estimate the policy parameters as well as the reward parameters.
3 Problem Formulation
[…]
Batched Contextual Bandit (BCB) Setting: Following the setups in linear contextual bandit literature [Dimakopoulou et al., 2019, Yang et al., 2021, Li et al., 2010], for any context $\bm{s}_{i}\in\mathcal{S}\subseteq\mathbb{R}^{d}$, we assume that the expectation of the observed reward $R_{i}$ from environment is determined by unknown true reward parameters $\bm{\theta}^{*}\in\mathbb{R}^{d}$: $\mathbb{E}[R_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{s}_{i}]=\langle\bm{\theta}^{*},\bm{s}_{i}\rangle.$ In this paper, our focus is on the generalized version of the traditional contextual bandit called the batched contextual bandit (BCB) setting, which has gained considerable attention in the field of bandit theory and applications [Han et al., 2020, Zhang et al., 2021]. For an expert policy in BCB setting, we define expert’s decision-making progress is partitioned into $N$ episodes, and in each episode, expert consists of two phases: (1) the online decision-making chooses the action (i.e., candidate context, and also behavior) for execution from each step’s candidate context set following the updated and fixed expert policy $f$ for $B$ steps ($B$ is also called batch size), and finally stores all context set & action pairs and the observed rewards of the executed actions (contexts) into the data buffer $\mathcal{D}$ (expert’s behavioral evolution history with rewards); (2) the policy updating approximates the optimal policy based on the received actions and rewards in $\mathcal{D}$. Note that for expert policy itself to evolve, expert can observe the reward of executed actions. Also, data buffer stores the indexes of the executed actions rather than actions themselves since each step’s context set contains all actions, including the executed actions. This saves the storage space for practical applications.
Figure 2 shows the content we can observe, where reward of executed actions (behaviors) is unaccessible. Then, we introduce the formal definition of inverse bandit problem with behavioral evolution history in BCB setting as follows.
Definition 1 (Inverse Bandit Problem with Behavioral Evolution History in BCB Setting).
Denote that $\bm{s}_{I}$, $I\in\mathcal{I}$, is a candidate context, =$\mathcal{I}=\{I_{j}\}_{j\in[M]}$ is the action’s index space containing M action indexes, $\mathcal{S}\subseteq\mathbb{R}^{d}$ is the context space whose dimension is $d$, $\mathcal{S}_{n,b}=\{\bm{s}_{I}\}_{I\in\mathcal{I}}\subseteq\mathcal{S}$ is the candidate context set for expert at step $b$ in the $n$-th episode, $I_{n,b}\in\mathcal{I}$ is the index of the expert policy’s executed action (behavior) at step $b$ in the $n$-th episode. Observe the behavioral evolution history generated by an expert with the format of $\mathcal{D}^{\mathrm{U}}=\bigcup_{n\in[N]}\mathcal{D}_{n}^{\mathrm{U}}$, where $\mathcal{D}_{n}^{\mathrm{U}}=\{(\mathcal{S}_{n,b},I_{n,b})\}_{b\in[B]}$, $\forall n\in[N]$, i.e., the reward feedbacks cannot be observed, $B$ is the batch size. We assume expert policy gets its parameters as $\{\bm{\theta}_{n}\}_{n\in[N]}$ through its online update method, and assume $\bm{\theta}_{N}$ is the closest estimation for reward parameter generated by expert policy in the evolution. We want to learn both expert policy parameter $\bm{\theta}_{N}$ and reward parameter $\bm{\theta}^{*}$ accurately through $\mathcal{D}^{\mathrm{U}}$.
4 IBCB: The Proposed Approach
In this section, we proposed Inverse Batched Contextual Bandit approach named IBCB which can efficiently learn from expert’s behavioral evolution history.
Policy Updating in BCB Setting: At the end of the $n$-th episode in the online learning period, the context vectors (corresponding to the executed actions) and their rewards are observed, and are stored into a context matrix $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ and a reward vector $\bm{R}_{n}\in\mathbb{R}^{B}$, respectively. In particular, each row index of $\bm{S}_{n}$ and $\bm{R}_{n}$ corresponds to the index of the step in which the context-reward pair was received. We introduce the updating process of the reward parameter vector $\bm{\theta}_{n}$ in BCB. We first concatenate the context and reward matrices from the previous episodes:
$\bm{L}_{n}=[\bm{S}_{0};\bm{S}_{1};\cdots;\bm{S}_{n}]\in\mathbb{R}^{nB\times d},$
$\bm{T}_{n}=[\bm{R}_{0};\bm{R}_{1};\cdots;\bm{R}_{n}]\in\mathbb{R}^{nB}.$
Then, the updated parameter vector $\bm{\theta}_{n+1}$ can be obtained by solving the following ridge regression: for $n=0,1,\ldots,N-1$,
$\bm{\theta}_{n+1}=\operatornamewithlimits{arg\,min}_{\bm{\theta}\in\mathbb{R}^{d}}\left\|\bm{L}_{n}\bm{\theta}-\bm{T}_{n}\right\|_{2}^{2}+\lambda\|\bm{\theta}\|_{2}^{2},$ (1)
where $\lambda>0$ is the regularization parameter. The closed least squares solution of Eq.(1) is used for estimating $\bm{\theta}_{n+1}$:
$\bm{\theta}_{n+1}=\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{b}_{n+1},$ (2)
where $\bm{\Phi}_{n+1}=\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n}=\sum_{i\in\left[n\right]}\bm{S}_{n}^{\top}\bm{S}_{n},\quad\bm{b}_{n+1}=\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n}=\sum_{i\in\left[n\right]}\bm{S}_{n}^{\top}\bm{R}_{n},\quad\bm{\Psi}_{n+1}=\lambda\bm{I}_{d}+\bm{\Phi}_{n+1}.$
Algorithm 1 Batched Policy Updating in the $n$-th episode in BCB Setting
0:  Policy $f_{n}$, data buffer $\mathcal{D}_{n}$, $\lambda\geq 0$, $\alpha\geq 0$, $\bm{\Phi}_{1}=\bm{O}_{d}$, $\bm{b}_{1}=\bm{0},$ batch size $B$, context space $\mathcal{S}$, action space $\mathcal{I}$
0:  Updated policy $f_{n+1}$
1:  Store the selected context vectors and the observed rewards into $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ and $\bm{R}_{n}\in\mathbb{R}^{B}$, respectively
2:  $\bm{\Phi}_{n+1}\leftarrow\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},$$\bm{b}_{n+1}\leftarrow\bm{b}_{n}+\bm{S}_{n}^{\top}\bm{R}_{n}$
3:  $\bm{\Psi}_{n+1}\leftarrow\lambda\bm{I}_{d}+\bm{\Phi}_{n+1},$$\bm{\theta}_{n+1}\leftarrow(\bm{\Psi}_{n+1})^{-1}\bm{b}_{n+1}$
[…]
5:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
[…]
7:  Draw $\tilde{\bm{\theta}}_{n+1}$ from $\mathcal{N}(\bm{\theta}_{n+1},\alpha^{2}(\bm{\Psi}_{n+1})^{-1})$
8:  $f_{n+1}$ selects the context as $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\tilde{\bm{\theta}}_{n+1},\bm{s}_{I}\rangle$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$
9:  Return $\bm{\theta}_{n+1},$  $\left(\bm{\Psi}_{n+1}\right)^{-1}$
Finally, as shown in Algorithm 1, the batched UCB policy updates the policy parameters and then selects the context $\bm{s}_{I_{n,b}}$ by applying the policy $f_{n+1}$ according to the following rule: $I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{I}]^{\frac{1}{2}}$, where $\bm{s}_{I}\in\mathcal{S}_{n,b}$. On the other hand, the batched Thompson sampling policy is a randomized policy that samples the policy parameters from a Gaussian distribution. However, as demonstrated in Theorem 1, we can prove that the batched Thompson sampling policy can also be expressed in the same regularized form as the UCB policy.
Theorem 1.
For batched Thompson sampling policy in batched policy updating (i.e., step 7 and 8 in Algorithm 1), we can obtain the reparameterized result of $f_{n+1}$ as:
$I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}]^{\frac{1}{2}}z$ (3)
where $z$ is a Gaussian random number drawn from $\mathcal{N}(0,1)$.
IBCB’s Formulation: We will illustrate the formulation using two actions as an example, which can be easily extended to accommodate multiple actions by adding additional constraints. Specifically, given $\mathcal{S}_{n+1,b}=\{\bm{s}_{1},\bm{s}_{2}\}$, =$\mathcal{I}=\{1,2\}$, if action 1 is executed at step $b$ in the $(n+1)$-th episode, we have
$\left\langle\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{b}_{n+1},\bm{s}_{1}\right\rangle+\alpha[\bm{s}_{1}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{1}]^{\frac{1}{2}}$
$\geq\left\langle\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{b}_{n+1},\bm{s}_{2}\right\rangle+\alpha[\bm{s}_{2}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{2}]^{\frac{1}{2}},$
which is equivalent to
$\left\langle\bm{b}_{n+1},\left(\bm{\Psi}_{n+1}\right)^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (4)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\left(\bm{\Psi}_{n+1}\right)^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\}.$
Setting $n=1$ in Eq.(4), from $\bm{b}_{2}=\bm{b}_{1}+\bm{S}_{1}^{\top}\bm{R}_{1}=\bm{S}_{1}^{\top}\bm{R}_{1}$ we get
$\left\langle\bm{R}_{1},\bm{S}_{1}\bm{\Psi}_{2}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (5)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\bm{\Psi}_{2}^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\bm{\Psi}_{2}^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\},$
where $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}_{2,b},\forall b\in[B].$
Similarly, setting $n=2$ in Eq.(4), from $\bm{b}_{3}=\bm{b}_{2}+\bm{S}_{2}^{\top}\bm{R}_{2}=\bm{S}_{1}^{\top}\bm{R}_{1}+\bm{S}_{2}^{\top}\bm{R}_{2}$ we have
$\left\langle\bm{S}_{1}^{\top}\bm{R}_{1}+\bm{S}_{2}^{\top}\bm{R}_{2},\bm{\Psi}_{3}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (6)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\bm{\Psi}_{3}^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\bm{\Psi}_{3}^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\},$
where $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}_{3,b},\forall b\in[B].$
Then, for all $n\in[N-1]$, given a candidate action set $\mathcal{S}_{n+1,b}=\{\bm{s}_{1},\bm{s}_{2}\}$, if action 1 is executed at step $b$ in the $(n+1)$-th episode (i.e., $\bm{s}_{1}$ is selected), we can obtain that, for $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}_{2,b},\forall b\in[B].$
$\sum_{i\in[n]}\left\langle\bm{R}_{i},\leavevmode\nobreak\ \bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (7)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\},$
otherwise action 2 is selected for execution at step $b$ in the $(n+1)$-th episode.
Taking expectations of both sides of the inequality Eq.(7), we have
$\mathbb{E}$ $\left[\sum_{i\in[n]}\left\langle\bm{R}_{i},\leavevmode\nobreak\ \bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle\bigg{|}\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{n}\right]$ (8)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\},$
Combining linear contextual bandit’s reward assumption in section 3, Eq.(8) is equivalent to
$\sum_{i\in[n]}\left\langle\bm{\theta}^{*},\leavevmode\nobreak\ \bm{S}_{i}^{\top}\bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (9)
$\geq\alpha\left\{[\bm{s}_{2}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{2}]^{\frac{1}{2}}-[\bm{s}_{1}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{1}]^{\frac{1}{2}}\right\}.$
Under the constraints Eq.(9), we can obtain the minimum-norm solution $\bm{\theta}^{*}$ by solving the following optimization problem with inequality constraints:
$\min\limits_{\bm{\theta}\in\mathbb{R}^{d}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{1}{2}\|\bm{\theta}\|_{2}^{2}$ (10)
s.t. $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i\in[n]}\left\langle\bm{\theta},\leavevmode\nobreak\ \bm{S}_{i}^{\top}\bm{S}_{i}\left[\bm{\phi}_{n}(\bm{s}_{\bar{I}_{n,b}})-\bm{\phi}_{n}(\bm{s}_{I_{n,b}})\right]\right\rangle$
$\leq\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right],\quad n\in[N],\leavevmode\nobreak\ b\in[B],$
where $\bm{\phi}_{n}(\bm{s})=\bm{\Psi}_{n+1}^{-1}\bm{s}$, $H_{n}(\bm{s})=\left[\bm{s}^{\top}\bm{\phi}_{n}(\bm{s})\right]^{\frac{1}{2}}$. Eq.(10) is the final formulation of IBCB, which proposed a simple quadratic programming (QP) problem with large scale linear constraints. Several tools have been introduced to efficiently solve QP problems [Gertz and Wright, 2003, Ferreau et al., 2014, Stellato et al., 2020]. Among these approaches, OSQP [Stellato et al., 2020] is the best to achieve almost linear complexity growth in problems’ dimension for large scale low-accuracy Random QP problems, which can be applied to our IBCB.
When we assume that original expert algorithm is SBTS, we can simply apply a random parameter for $H_{n}(\bm{s})$ as $H_{n}(\bm{s})=\left[\bm{s}^{\top}\bm{\phi}_{n}(\bm{s})\right]^{\frac{1}{2}}z$, where $z\sim\mathcal{N}(0,1)$ by following Theorem 1 and clear $H_{n}(\bm{s})$ to 0 by following Eq.(8)’s expectation setting. This indicates that IBCB is a unified framework for both deterministic and randomized BCB.
Existing inverse bandit algorithm Hüyük et al. [2022] estimates parameters using an EM-like algorithm, which iteratively maximizes the likelihood estimation. However, this algorithm suffers from low computational efficiency as it requires a large number of iterations and samples to achieve accurate estimation. Furthermore, the assumption of expert updates after each action under the fully-online setting leads to a rapid increase in computation time when sampling a large number of actions, making it unsuitable for the BCB setting. In contrast, the proposed IBCB capitalizes on the BCB assumption, which provides a unified framework for both deterministic and randomized bandit policies. By formulating the inverse bandit problem as a quadratic optimization problem with a significant number of linear constraints, IBCB offers a more efficient solution through leveraging the capabilities of the OSQP optimizer.
5 Experiments Settings
In this section, we evaluate the effectiveness and efficiency of IBCB. Detailed training settings of experts, all baselines and IBCB can be found in Appendix B.
5.1 Baselines
IBCB was compared with several existing original expert algorithms and imitation learning (IL) algorithms, including:
Original Expert Algorithms with Rewards (SBUCB & SBTS): We used SBUCB and SBTS for original expert algorithms (served as backbone algorithms in recommendations). These algorithms’ performances are the oracle limit above all of other IL algorithms and IBCB since they have access to the rewards. SBUCB is a batched version of LinUCB [Li et al., 2010], which updates the policy after receiving a batch of feedback data [Han et al., 2020]. SBTS is a batched version of Linear Thompson Sampling (LinTS) [Dimakopoulou et al., 2019].
Behavior Cloning (BC) Algorithms: We used LinearSVC [Pedregosa et al., 2011] model (namely BaseSVC) for BC baseline, since the we assumed reward enviornment is linear in Section 3.
Inverse Reinforcement Learning (IRL) Algorithms: We used Bayesian Inversive Reinforcement Learning (B-IRL) [Ramachandran and Amir, 2007], Bayesian Inversive Contextual Bandits (B-ICB) [Hüyük et al., 2022] to serve as IRL baselines.
5.2 Datasets
We designed synthetic and real-world dataset for experiments. For both of them, we split the whole dataset into two subsets: the Online Learning (the OL-data) phase, and the Batch Testing (the BT-data) phase. Note that expert only update its policy parameters in OL-data, and do not update parameter in BT-data phase so we can design experiments on estimating expert policy’s parameters on BT-data.
Synthetic Dataset. We generated the synthetic data for both OL-data and BT-data as follows: numbers of episodes $N=20$, batch size $B=1000$, number of candidate contexts (actions) is $M=10$ at any single step of an episode. Other details can be found in Appendix B.1.
Real-World Dataset. We used MovieLens 100K  [^1] https://grouplens.org/datasets/movielens/100k/ (ML-100K) dataset as our real-world experiment dataset. For OL-data, we set numbers of episodes $N=20$, batch size $B=1000$, and for BT-data, we set $N=50$, $B=1000$. Number of candidate contexts (actions) is $M=10$ at any single step of an episode in both OL-data and BT-data phase. Other details can be found in Appendix B.1.
5.3 Evaluation Protocol
Online Train Log Fitness (OL-Fitness) is used to evaluate the similarity between the parameters learned by various algorithms and the parameters of the true reward environment. Specifically, the learned parameters is used to fit the reward environment, and then the expert interacts with it to obtain contexts and rewards for online learning. The resulting series of recommended (executed) actions from this process are then compared to the actions obtained by the expert interacting with the true reward environment (OL-data environment) as $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}\mathbb{I}(I_{n,b-\mathrm{learned}}=I_{n,b-\mathrm{reward}})$, where $I_{n,b-\mathrm{learned}}$ is the index of action which expert online-recommended on estimated parameters learned by IL or IBCB at step $b$ of $n$-th episode in OL-data phase, and $I_{n,b-\mathrm{reward}}$ is the index of action which expert online-recommended on true reward parameters at step $b$ of $n$-th episode in OL-data phase. Higher OL-Fitness means more similarity the parameters learned by an algorithm has for the reward parameters.
Batch Test Log Fitness (BT-Fitness) is used to evaluate the the similarity between the parameters/models learned by various algorithms and the final expert policy’s parameters. BT-Fitness compares the actions learned parameters/models recommended to the actions final experts’ policy recommended in BT-data phase as $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}\mathbb{I}(I_{n,b-\mathrm{learned}}=I_{n,b-\mathrm{policy}})$, where $I_{n,b-\mathrm{learned}}$ is the index of action recommended by IL models or IBCB at step $b$ of $n$-th episode in BT-data phase, and $I_{n,b-\mathrm{policy}}$ is the index of action recommended by final experts’ policy at step $b$ of $n$-th episode in BT-data phase. The higher BT-Fitness an algorithm achieves, the more similarity it has for the final expert policy’s parameters.
Batch Test Average Reward (BT-AR) is used to evaluate the reward performance of the parameters/models learned by various algorithms for the BT-data phase, based on the average reward feedback obtained after the recommended actions interact with the BT-data reward environment as $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}R_{n,b},\ R_{n,b}\in\{0,1\}$, where $R_{n,b}$ is the recommended action’s observed reward at step $b$ of $n$-th episode in BT-data phase. The higher the average reward, the more competitive the recommended actions are in the BT-data reward environment (may even be an out-of-distribution data environment), and the better reward performance an algorithm has.
6 Experimental Results and Analyses
We compared several indicators between IBCB with other baselines in this section. We used data sampling method to train B-ICB (namely B-ICB-S, where ‘-S’ means ‘used sampled data’) since B-ICB has not yet been adapted to batched version, so B-ICB can only be trained in environment that quantity of train data is not so large.
6.1 Synthetic Experimental Results & Discussions
Table 1: Reward parameters’ estimation, where $\mathrm{std}$ denotes the standard deviation of zero-mean Gaussian noise added to the final reward.
Algorithm Online Train Log Fitness (Original Expert is SBUCB)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019
B-IRL 0.827±0.012 0.808±0.039 0.806±0.037 0.802±0.036
B-ICB-S 0.805±0.020 0.771±0.031 0.780±0.034 0.771±0.024
IBCB (Ours) 0.882±0.012 0.870±0.028 0.871±0.033 0.865±0.029
Algorithm Online Train Log Fitness(Original Expert is SBTS)
std=0 std=0.03 std=0.07 std=0.1
SBTS 0.857±0.017 0.876±0.012 0.872±0.014 0.871±0.008
B-IRL 0.457±0.020 0.468±0.029 0.480±0.023 0.476±0.039
B-ICB-S 0.424±0.020 0.451±0.036 0.457±0.016 0.454±0.024
IBCB (Ours) 0.567±0.031 0.607±0.030 0.622±0.026 0.613±0.039
Table 2: Train time comparison. Experts (SBUCB & SBTS) do not need to train since they produce the train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm Train Time (sec.) (Original Expert is SBUCB)
std=0 std=0.03 std=0.07 std=0.1
SBUCB / / / /
B-IRL 36.997±0.803 37.67±1.328 39.91±3.419 37.424±0.563
BaseSVC 116.86±20.63 116.39±26.58 132.18±19.86 117.25±19.47
B-ICB-S 86.365±0.827 87.599±4.489 86.85±1.162 85.713±0.676
IBCB (Ours) 6.622±2.621 7.992±3.218 7.186±1.577 8.615±4.599
Algorithm Train Time (sec.) (Original Expert is SBTS)
std=0 std=0.03 std=0.07 std=0.1
SBTS / / / /
B-IRL 38.46±3.550 36.419±0.187 37.843±3.486 43.469±6.228
BaseSVC 96.603±9.619 87.310±8.435 89.240±6.566 90.472±8.494
B-ICB-S 91.166±1.164 89.193±1.477 89.892±1.162 89.483±1.585
IBCB (Ours) 8.236±2.153 5.226±2.474 7.476±4.323 9.415±2.835
Table 3: Expert policy (SBUCB) parameters’ estimation. Expert policy produce same actions compared with itself, so Batch Test Log Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness
std=0 std=0.03 std=0.07 std=0.1
SBUCB 1 1 1 1
B-IRL 0.828±0.005 0.820±0.013 0.820±0.014 0.818±0.016
BaseSVC 0.791±0.008 0.781±0.020 0.781±0.020 0.777±0.020
B-ICB-S 0.800±0.032 0.767±0.055 0.780±0.045 0.776±0.022
IBCB (Ours) 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.641±0.01 0.629±0.028 0.630±0.024 0.622±0.024
B-IRL 0.588±0.015 0.571±0.037 0.572±0.034 0.561±0.034
BaseSVC 0.577±0.017 0.559±0.042 0.559±0.037 0.547±0.037
B-ICB-S 0.581±0.006 0.553±0.024 0.560±0.025 0.552±0.025
IBCB (Ours) 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
Table 4: Expert policy’s (SBTS) parameters estimation
Algorithm Batch Test Log Fitness
std=0 std=0.03 std=0.07 std=0.1
SBTS 1 1 1 1
B-IRL 0.816±0.007 0.82±0.015 0.816±0.008 0.816±0.006
BaseSVC 0.789±0.011 0.807±0.009 0.801±0.015 0.798±0.012
B-ICB-S 0.740±0.059 0.742±0.089 0.741±0.074 0.752±0.076
IBCB (Ours) 0.856±0.019 0.877±0.011 0.873±0.018 0.867±0.018
Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.1
SBTS 0.650±0.020 0.662±0.025 0.661±0.025 0.656±0.024
B-IRL 0.595±0.026 0.615±0.028 0.611±0.031 0.605±0.030
BaseSVC 0.586±0.032 0.611±0.035 0.607±0.037 0.600±0.034
B-ICB-S 0.571±0.021 0.589±0.022 0.588±0.021 0.584±0.023
IBCB (Ours) 0.609±0.031 0.631±0.033 0.628±0.035 0.621±0.034
Table 1 to Table 2 report the OL-Fitness and training time of experts (SBUCB & SBTS, oracle limit), baselines, and the IBCB. We observed that IBCB achieved higher OL-Fitness, and its training time significantly outperformed other baselines. The assumption and conformity of the BCB setting enable IBCB to achieve the best performance in reward parameter estimation. Thanks to the OSQP solver, IBCB can solve large-scale low-accuracy QP problems within almost linear complexity growth of the problem dimension, which leads to significantly faster training than other baselines. BaseSVC’s iteration design enjoys a complexity growth between linear and quadratic but close to quadratic, resulting in slower training than IBCB. B-IRL has a linear complexity growth but needs a large quantity of Metropolis-Hastings sampling iterations to achieve better performance. We analyze B-ICB’s complexity as quadratic due to its assumption that the expert updates after every single step and its Bayesian updating method derived from B-IRL. Therefore, it is not surprising that B-ICB achieves a very slow training speed even using sampled data.
Table 3-Table 4 reports BT-Fitness and BT-AR of different algorithms. We observed that IBCB achieved higher BT-Fitness and BT-AR, demonstrating the effectiveness of our reward imputation in Eq.(9) for behavioral evolution history under unaccessible feedback generated by expert, which also confirmed that IBCB can learn expert policy’s parameter with great effectiveness and efficiency. Similar with the results in former paragraph, IBCB benefits a lot from the assumption and conformation to BCB setting. Even in the OOD scenario (see Table 10-Table 11 in Appendix B.4.1, IBCB stills outperforms other baselines, which indicates its robustness for data with shifted distribution.
6.2 Real-World ML-100K Experimental Results & Discussions
Table 5: Reward parameters’ estimation & train time comparison. Experts do not need to train since they produce train logs (behavioral evolution history logs) for baselines and IBCB. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison. Original Expert in Subtable 5(a) is SBUCB, and Original Expert in Subtable 5(b) is SBTS.
Algorithm OL-Fitness Train Time (sec.)
SBUCB 0.891±0.006 /
B-IRL 0.830±0.010 43.669±1.282
BaseSVC / 162.767±12.135
B-ICB-S 0.746±0.032 147.474±4.633
IBCB (Ours) 0.870±0.014 15.023±3.854
(a) Original Expert Algorithm is SBUCB
Algorithm OL-Fitness Train Time (sec.)
SBTS 0.868±0.015 /
B-IRL 0.851±0.015 47.005±4.311
BaseSVC / 192.83±33.199
B-ICB-S 0.721±0.030 145.441±5.916
IBCB (Ours) 0.860±0.026 13.046±3.76
(b) Original Expert Algorithm is SBTS
Table 6: Expert policy parameters’ estimation. Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. Original Expert in Subtable 6(a) is SBUCB, and Original Expert in Subtable 6(b) is SBTS.
Algorithm BT-Fitness BT-AR
SBUCB 1 0.193±0.005
B-IRL 0.864±0.008 0.189±0.005
BaseSVC 0.804±0.012 0.189±0.005
B-ICB-S 0.767±0.028 0.187±0.007
IBCB (Ours) 0.884±0.021 0.192±0.006
(a) Original Expert Algorithm is SBUCB
Algorithm BT-Fitness BT-AR
SBTS 1 0.200±0.005
B-IRL 0.870±0.030 0.199±0.007
BaseSVC 0.801±0.019 0.198±0.007
B-ICB-S 0.759±0.035 0.196±0.004
IBCB (Ours) 0.844±0.026 0.199±0.005
(b) Original Expert Algorithm is SBTS
Results on ML-100K dataset of different indicators are reported in Table 5-Table 6. Similar to the results on synthetic dataset, IBCB achieved higher performance than other baselines in several indicators, despite the BT-Fitness on SBTS backbone. Since dimension of real-world’s context vector is very large and embedding of the context vector is very sensitive to randomness, we assumed this phenomenon occurred due to that randomness in real-world would undermine IBCB’s performance because IBCB simply sets right side of Eq.10 to be 0 and ignores the randomness. From the outstanding results of training speed and average reward, we concluded that IBCB can also adapt to learning real-world behavioral evolution history generated by expert with great efficiency.
6.3 Ablation Study
In this subsection, we discussed several extreme situations (mentioned in Introduction 1) that imitation learning (IL) algorithms and IBCB may face with. Training logs (expert’s behavioral evolution history) may update in a delayed manner in online learning, which means that when training models, we may only have access to former part of expert’s behavioral evolution history data (e.g. IL and IBCB can only obtained the former half part of the data, but original expert has evolved with the whole training data), and contradictory data may appear. Since that original expert policy’s hyper parameter ($\alpha$ in Algorithm 1, detailed settings shown in Appendix B.2.1) is also unaccessible for IBCB, we also checked the influence of IBCB’s hyper parameter’s ($\alpha$ in Eq.(10) ) setting.
Detailed results and analyses of IBCB’s hyper parameter test on synthetic dataset and the whole ablation study on ML-100K dataset can be found in Appendix B.4.3.
Table 7: Reward parameters’ estimation on synthetic dataset (using first half of expert’s behavior evolution history logs). BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison.
Algorithm Online Train Log Fitness (Half Train)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019
B-IRL 0.752±0.018 0.729±0.051 0.728±0.050 0.721±0.047
B-ICB-S 0.730±0.039 0.727±0.045 0.721±0.063 0.717±0.041
IBCB (Ours) 0.842±0.016 0.827±0.037 0.825±0.041 0.819±0.038
Table 8: Expert policy parameters’ estimation on synthetic dataset (using first half of expert’s behavioral evolution history logs). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. Original Expert is SBUCB.
Algorithm Batch Test Log Fitness (Half Train)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 1 1 1 1
B-IRL 0.718±0.012 0.699±0.031 0.702±0.035 0.694±0.037
BaseSVC 0.656±0.019 0.638±0.039 0.638±0.043 0.627±0.045
B-ICB-S 0.695±0.054 0.707±0.069 0.695±0.084 0.688±0.043
IBCB (Ours) 0.775±0.014 0.762±0.035 0.761±0.041 0.757±0.038
Algorithm Batch Test Average Reward (Half Train)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.641±0.010 0.629±0.028 0.630±0.024 0.622±0.024
B-IRL 0.545±0.020 0.522±0.047 0.522±0.046 0.510±0.046
BaseSVC 0.520±0.023 0.497±0.053 0.497±0.052 0.482±0.051
B-ICB-S 0.549±0.019 0.538±0.024 0.540±0.030 0.529±0.033
IBCB (Ours) 0.569±0.019 0.549±0.048 0.548±0.046 0.537±0.045
Table 9: Contradictory (duplicated) data (appeared at expert’s OL-data behavioral evolution history logs) on synthetic dataset testing. $\mathrm{dup}$ denotes the quantity of duplicated episodes in OL-data phase. OL-Fitness and Train Log Fitness (BC) results were shown in Appendix B.4.2. Original Expert is SBUCB.
Algorithm Batch Test Log Fitness ($\mathrm{dup}<5$)
dup=0 dup=1 dup=2 dup=3 dup=4
SBUCB 1 1 1 1 1
B-IRL 0.820±0.013 0.820±0.012 0.820±0.013 0.820±0.013 0.819±0.014
BaseSVC 0.781±0.020 0.783±0.019 0.782±0.020 0.781±0.018 0.780±0.021
IBCB (Ours) 0.843±0.028 0.843±0.027 0.842±0.028 0.843±0.027 0.843±0.029
Algorithm Batch Test Log Fitness ($\mathrm{dup}\geq 5$)
dup=5 dup=6 dup=7 dup=8 dup=9
SBUCB 1 1 1 1 1
B-IRL 0.818±0.014 0.819±0.014 0.818±0.011 0.818±0.011 0.819±0.014
BaseSVC 0.780±0.020 0.779±0.020 0.778±0.018 0.776±0.018 0.776±0.021
IBCB (Ours) 0.841±0.029 0.843±0.029 0.842±0.029 0.841±0.027 0.842±0.03
Table 7–Table 9 reports the results of different ablation studies (exclude IBCB’s hyper parameter test, see details in Appendix B.4.3) on synthetic dataset. We have the following observations: (1) When all IL baselines and IBCB only have the access to the first half of expert’s behavioral evolution history logs, i.e., novice expert’s evolution history, IBCB can have better performance, compared with the degree of decline in all algorithms from Table 1 and Table 3. (2) Contradictory data undermines BaseSVC’s (BC model) fitness on expert’s behavior evolution history and BT-Fitness, but models which learns reward parameters, including IBCB, are not influenced by this phenomenon, which demonstrate IBCB’s robustness for extreme situations.
7 Conclusion
This paper aims to address the problem of efficiently learning behaviours from the expert’s evolution history. Specifically, we proposed an inverse batched contextual bandit model called IBCB. The proposed IBCB gives a unified framework for both deterministic and randomized bandit policies, and solves the problem of learning from evolution history with unaccessible rewards through a simple quadratic programming problem. Experimental results demonstrated the effectiveness and efficient training of IBCB on both synthetic and real-world scenarios. Besides, IBCB can also be generalized to out-of-distribution and contradictory data scenarios with great robustness.
Appendix A Appendix: Detailed Proofs
A.1 Experiment Codes
Codes are available in the following URL: https://anonymous.4open.science/r/IBCB-F25F.
A.2 Proof of Theorem 1
Proof.
Proof of the reparameterization of the SBTS algorithm
From Algorithm 1 we get the original form of SBTS’s parameter:
$\tilde{\bm{\theta}}_{n+1}\sim\mathcal{N}(\bm{\theta}_{n+1},\alpha^{2}(\bm{\Psi}_{n+1})^{-1})$ (11)
Since $\bm{\Psi}_{n+1}^{-1}$ is symmetric and positive-definite, so we can reparameterize $\tilde{\bm{\theta}}_{n+1}$ as follows:
$\tilde{\bm{\theta}}_{n+1}=\bm{\theta}_{n+1}+\alpha\bm{B}_{n+1}\bm{z}$ (12)
where $\bm{B}_{n+1}\bm{B}_{n+1}^{\top}=\bm{\Psi}_{n+1}^{-1}$, and $\bm{z}\sim\mathcal{N}(0,\bm{I}_{d})$.
Then we can rewrite the step 5 of Algorithm 1 as:
$I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{z}$ (13)
Note that for any matrix $\bm{A}$ with dimension of $r\times p$, $\bm{A}$ with dimension of $s\times q$, any random variable $x$ with dimension of $p\times 1$, $y$ with dimension of $q\times 1$ and any constant vector $b$ with dimension of $r\times 1$, $d$ with dimension of $s\times 1$, where $r,p,s,q$ is constant number, we have:
$Cov(\bm{A}_{r\times p}\bm{x}_{p\times 1}+\bm{b}_{r\times 1},\bm{C}_{s\times q}\bm{y}_{q\times 1}+\bm{d}_{s\times 1})$ (14)
$=\bm{A}Cov(\bm{x},\bm{y})\bm{C}^{\top}$
where $Cov(\cdot,\cdot)$ means the covariance matrix.
Combining $\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{z}$ and (14) yields that:
$Cov(\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{z},\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{z})$ $=\bm{s}_{I}^{\top}\bm{B}_{n+1}Cov(\bm{z},\bm{z})\bm{B}_{n+1}^{\top}\bm{s}_{I}$ (15)
$=\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{B}_{n+1}^{\top}\bm{s}_{I}$
$=\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}$
And $\bm{z}\sim N(0,\bm{I}_{d})$, so we have:
$\bm{s}_{I}^{\top}\bm{B}_{n+1}\bm{z}\sim\mathcal{N}(0,\|\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}\|^{2})$ (16)
Note that:
$\left[\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}\right]^{\frac{1}{2}}z\sim\mathcal{N}(0,\|\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}\|^{2})$ (17)
where $z\sim\mathcal{N}(0,1)$
So finally, after combining (16) and (17), we can rewritten (13) as:
$I_{n,b}\leftarrow\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\bm{\theta}_{n+1},\bm{s}_{I}\rangle+\alpha[\bm{s}_{I}^{\top}\bm{\Psi}_{n+1}^{-1}\bm{s}_{I}]^{\frac{1}{2}}z$ (18)
Appendix B Appendix: Experiment Details
B.1 Details of Dataset Settings
Synthetic Dataset: We generated the candidate context $\bm{s}_{I}$ with context dimension $d=10$ as a single context (behavior) for any step of an episode. At step $b$ of $n$-th episode in OL-data phase, candidate context set $\mathcal{S}_{n,b}\subseteq{\mathbb{R}^{d}}$ was drawn from a Gaussian distribution $\mathcal{N}(\mu_{s}\bm{1}_{d},\sigma_{s}^{2}\bm{I}_{d})$, where the means of candidate contexts were $\mu_{s}\in\left[1:-0.4:-2.6\right]$, and the standard deviation was $\sigma_{s}^{2}={0.05}^{2}$. To simulate out-of-distribution(OOD) scenario, the mean of the first candidate context in the BT-data phase was set to $1.4$. We set the observed reward of a given candidate context $\bm{s}_{I_{n,b}}\in\mathcal{S}_{n,b}$ at step $b$ in $n$-th episode as a sigmoid function $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)$, where each element of $\bm{w}_{\mathrm{reward}}\in\mathbb{R}^{d}$ was sampled from a Gaussian distribution as $\mathcal{N}(0.1,{0.01}^{2})$. To test the robustness of the model and simulate random selections of actions made by users, we also added a zero-mean Gaussian noise $\mathrm{std}\sim\mathcal{N}(0,\epsilon^{2}),\epsilon\in\{0.03,0.07,0.1\}$ in the reward, so final reward was set to be $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)+\mathrm{std}$. We also made adjustments for the ablation study period. Contradictory data was introduced as duplicated data: we cloned the first k episodes’ candidate context sets for the second k episode’s candidate context sets, and design k with the name of duplicated quantity $\mathrm{dup}\in\left[0:1:9\right]$. Total number of conflict executed actions (e.g. expert select action 1 at step $b$ of $n$-th episode, but select action 2 at step $b$ of $(n+k)$-th episode.) of the contradictory data was ascending when we increased $\mathrm{dup}$. We also controlled the training rate for all baselines and IBCB: we set the end episode of the exposed expert’s behavioral evolution history logs with constraint end rate $\mathrm{ce\_rate}\in\left[0,1\right]$ ($e.g.$ if $\mathrm{ce\_rate}=0.5$. we use the first $\left\lfloor 0.5N\right\rfloor$ episodes for baselines and IBCB training).
Real-World ML-100K Dataset: We used MovieLens 100K (ML-100K) dataset as our real-world data experiment dataset, which collected 5-star movie ratings from a movie recommendation service, including 943 unique users and 1682 unique movies(items). ML-100K dataset also provided various side information of both users and items. We split ML-100K dataset into OL-Data and BT-data according to the timestamps, where first half was treated as OL-data and the last half was BT-data. For reward setting, we treated the 4-5 star ratings as positive feedback(labeled with 1), and 1-3 star rating as negative feedback (labeled with 0). Through feature embedding [Wang et al., 2019] and principal component analysis (PCA), dimension of each [user, item] pair (candidate context, i.e., behavior) was reduced to 50, $i.e.,\ d=50$. For the candidate context set, we retained the original item for each interaction and then randomly sampled $9$ extra items from the entire item set (exclude the original item), so the number of candidates is $M=10$. In real recommendation, user-item interaction is sparse, meaning that not all of the items can be shown to the user and have feedback [Zhang et al., 2021, 2022]. To overcome this issue, we trained two Matrix Factorization (MF) [Koren et al., 2009] models using OL-data (denotes MF-OL) and BT-data (denotes MF-BT) respectively for different baselines and IBCB model to serve as the simulated real environment. AUCs of the two MF models were both over $90\%$, which assures that the simulated environment can provide nearly realistic feedbacks of users. At step $b$ of $n$-th episode of both OL-data and BT-data phase, the simulated environment (MF-OL/MF-BT) received a selected candidate context $\bm{s}_{I_{n,b}}$ and a context set $\mathcal{S}_{n,b}$ consisted of all candidate contexts in this step, getting all candidate context’s reward and sort out the candidate context $\bm{s}_{I_{\mathrm{max}}}$ with max reward $r_{I_{\mathrm{max}}}$. If selected candidate context $\bm{s}_{I_{n,b}}$ do not have the biggest reward compared to $\bm{s}_{I_{\mathrm{max}}}$, $i.e.,r_{I_{n,b}}<r_{I_{\mathrm{max}}}$, reward feedback would be directly set as 0, otherwise would be set as $r_{I_{\mathrm{max}}}$. If $r_{I_{\mathrm{max}}}>0.5$, then the final reward feedback would be 1, otherwise would be 0. Adjustments for the ablation study period was also made similar with the synthetic dataset. Since duplicated data will not appeared in a large scale in real world, we just controlled the training rate similarly as that in synthetic dataset’s setting.
Note that MF-OL is quite different compared to MF-BT (i.e., BT-data’s distribution is different to OL-data’s distribution), so experiments made on BT-data can directly provide the results of model’s generalization ability in out-of-distribution data.
B.2 Training Details
We have run all of the experiments on the server equipped with 16 Intel Xeon E5-2630 v4@2.20GHz cores, and during training, all of the baselines and IBCB only uses the CPU. Note that all baselines and IBCB were training on expert behavioral evolution history of Online Learning. Detailed settings of baselines and IBCB are shown as follows.
B.2.1 Original Expert Policy Settings
SBUCB: We set the hyper parameter ($\alpha$ in step 5 of Algorithm 1, i.e., exploration parameter) of SBUCB to 0.4 for both synthetic and real-world ML-100K dataset experiments. Note that this parameter is only effective in Online Learning phase since expert (SBUCB) do not update its policy parameters (i.e., explore) in Batch Test phase.
SBTS: We set the hyper parameter ($\alpha$ in step 7 of Algorithm 1, i.e., exploration parameter) of SBTS to 0.4 for synthetic dataset experiments and 0.05 for real-world ML-100K dataset experiments. Note that this parameter is only effective in Online Learning phase since expert (SBTS) do not update its policy parameters (i.e., explore) in Batch Test phase.
B.2.2 Baseline Settings
BaseSVC: We followed the setting recommended in $\mathrm{sklearn.svm.LinearSVC}$ and set the $\mathrm{max\_iteration}$ (i.e., the maximum number of iterations) to 1,000,000 to ensure that the model can converge in the end.
B-IRL: Following [Hüyük et al., 2022], we have run Metropolis-Hastings algorithm for 10,000 iterations to obtain 1,000 samples on synthetic dataset (500 samples and 5,000 iterations for real-world ML-100K dataset) with intervals of 10 iterations between each sample after 10,000 burn-in iterations (5,000 for real-world ML-100K dataset). The final estimated parameters are formed by simply averaging all 1,000 samples. (500 for real-world ML-100K dataset).
B-ICB: We followed the setting used in [Hüyük et al., 2022]’s work. We changed the setting of $\mathrm{compute\_rhox}$ period (i.e., computing the reward parameters) in real-world ML-100K dataset experiments. Since that ML-100K’s user embedding used one-hot encoding, which may make the final matrix $\bm{\beta}_{n}$ to concatenate in $\mathrm{compute\_rhox}$ be non-invertible, we replace the original $\mathrm{np.linalg.inv}$ method with $\mathrm{np.linalg.pinv}$ method to compute the pseudo inverse of $\bm{\beta}_{n}$, making sure experiment can be done without computational errors. And due to that $\bm{\beta}_{n}$ may be non-invertible, we also dismiss the $\mathrm{det}(\beta_{n}/\sigma^{2})$ during iteration, where $\mathrm{det}(\bm{A})$ means the determinant of matrix $\bm{A}$.
Data Sampling Rate for B-ICB denotes the proportion of training data B-ICB used for training since B-ICB cannot be run with large-scale training data. When the data sampling rate is $k\%$, we uniformly randomly sample $k\%$ of the original expert’s training data as the training data (behavioral evolution history) for B-ICB. Data sampling rate was set to be 1% to make sure that B-ICB can be trained in 100s on synthetic dataset experiments, and 0.5% to make sure that B-ICB can be trained in 150s on real-world ML-100K dataset experiments. Also, in ablation study section, we used this setting for B-ICB.
B.2.3 IBCB Settings
We used OSQP solver to solve Eq.(10)’s quadratic problem(QP) as mentioned. Tolerance levels of OSQP was set to be $\epsilon_{abs}=8\times 10^{-2},\epsilon_{rel}=8\times 10^{-2}$. If QP with the former setting cannot convergent, then setting would be modified as $\epsilon_{abs}=8\times 10^{-1},\epsilon_{rel}=8\times 10^{-1}$. For real-world ML-100K experiments with SBTS backbone, setting was $\epsilon_{abs}=3\times 10^{-1},\epsilon_{rel}=3\times 10^{-1}$. Also, when expert’s policy is SBTS, right-hand side of Eq.(10) should be set to a constant $\epsilon$ slightly greater than 0 to prevent the parameters learned by IBCB from being too close to the origin ($\bm{0}_{d}$). Therefore, we set IBCB’s $\epsilon$ to 0.01 in the experimental settings of the synthetic dataset and 0.001 in real-world ML-100K dataset when we assume expert’s policy is SBTS.
B.3 Details of the Evaluation Protocol
OL-Fitness: OL-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds. Note that BC algorithms cannot obtain reward parameters, and thus were not included in this comparison scope. But in dup-test period, we used ‘train log behaviors fitness’ to evaluate whether BC algorithm suffers from contradictory data.
BT-Fitness: BT-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds, and expert do not update its policy parameters in BT-data phase.
B.4 Results Details
All results were averaged over 5 different runs, and sampled the data over 5 different random seeds for each train data log (behavioral evolution history log) of expert on both synthetic and real-world ML-100K dataset.
B.4.1 Results of Out-of-distribution Data on Synthetic Dataset
Table 10: Expert policy (SBUCB) parameters’ estimation (Out-of-distribution data). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 1 1 1 1
B-IRL 0.802±0.005 0.793±0.013 0.795±0.015 0.793±0.016
BaseSVC 0.764±0.009 0.756±0.021 0.756±0.022 0.750±0.023
B-ICB-S 0.775±0.039 0.738±0.065 0.755±0.050 0.754±0.025
IBCB (Ours) 0.834±0.012 0.822±0.031 0.821±0.033 0.816±0.029
Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.709±0.015 0.695±0.036 0.694±0.031 0.684±0.030
B-IRL 0.642±0.019 0.622±0.047 0.622±0.042 0.608±0.041
BaseSVC 0.627±0.021 0.607±0.052 0.606±0.046 0.591±0.046
B-ICB-S 0.632±0.008 0.600±0.030 0.607±0.032 0.596±0.032
IBCB (Ours) 0.654±0.021 0.632±0.054 0.632±0.048 0.618±0.046
Table 11: Expert policy (SBTS) parameters’ estimation (Out-of-distribution data). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBTS 1 1 1 1
B-IRL 0.792±0.005 0.795±0.010 0.793±0.006 0.793±0.005
BaseSVC 0.767±0.015 0.782±0.017 0.780±0.019 0.776±0.014
B-ICB-S 0.714±0.065 0.710±0.096 0.712±0.081 0.724±0.086
IBCB (Ours) 0.839±0.022 0.859±0.021 0.858±0.022 0.852±0.022
Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBTS 0.719±0.028 0.739±0.032 0.735±0.032 0.729±0.032
B-IRL 0.650±0.034 0.676±0.037 0.671±0.039 0.663±0.037
BaseSVC 0.640±0.041 0.670±0.045 0.666±0.047 0.658±0.044
B-ICB-S 0.620±0.028 0.643±0.028 0.642±0.026 0.637±0.027
IBCB (Ours) 0.667±0.040 0.697±0.043 0.693±0.044 0.684±0.044
As shown in Table 10 and Table 11, IBCB still outperforms other IL baselines in expert policy’s parameter estimation on Out-of-distribution data.
B.4.2 Detailed Results of Contradictory Data Testing on Synthetic Dataset
Table 12: Contradictory (duplicated) data (appeared at expert’s OL-data behavioral evolution history logs) on synthetic dataset testing. $\mathrm{dup}$ denotes the quantity of duplicated episodes in OL-data phase.
algorithm Online Train Log Fitness ($\mathrm{dup}<5$)
dup=0 dup=1 dup=2 dup=3 dup=4
SBUCB 0.879±0.016 0.880±0.015 0.880±0.015 0.879±0.015 0.880±0.018
B-IRL 0.808±0.039 0.808±0.037 0.808±0.038 0.806±0.038 0.807±0.038
IBCB (Ours) 0.870±0.028 0.870±0.03 0.868±0.029 0.867±0.030 0.869±0.030
Algorithm Train Log Behaviors Fitness (BC) ($\mathrm{dup}<5$)
dup=0 dup=1 dup=2 dup=3 dup=4
BaseSVC
0.823±0.026 0.825±0.025 0.824±0.026 0.821±0.026 0.820±0.026
Algorithm Online Train Log Fitness ($\mathrm{dup}\geq 5$)
dup=5 dup=6 dup=7 dup=8 dup=9
SBUCB 0.879±0.015 0.880±0.016 0.879±0.015 0.879±0.015 0.878±0.017
B-IRL 0.807±0.037 0.809±0.037 0.810±0.037 0.809±0.037 0.808±0.037
IBCB (Ours) 0.868±0.029 0.870±0.030 0.872±0.029 0.870±0.028 0.869±0.029
Algorithm Train Log Behaviors Fitness (BC) ($\mathrm{dup}\geq 5$)
dup=5 dup=6 dup=7 dup=8 dup=9
BaseSVC
0.819±0.025 0.819±0.025 0.819±0.024 0.817±0.023 0.816±0.023
Note that BaseSVC cannot obtain reward parameters, and thus were not included in OL-Fitness. We used ‘Train Log Behaviors Fitness’ to evaluate whether BaseSVC suffers from contradictory data and make comparison with IBCB and other baselines. As shown in Table 12, contradictory data (same context in different time periods) undermines BaseSVC’s train log behaviors fitness, but IBCB stayed stable in OL-Fitness.
B.4.3 Results of IBCB’s Parameter Test on Synthetic Dataset
Table 13: Reward parameters’ estimation and training on synthetic dataset (IBCB’s parameter test). Experts do not need train since they produce train logs (behavioral evolution history logs) for baselines and IBCB. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison. Original Expert is SBUCB.
Algorithm Online Train Log Fitness (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019
B-IRL 0.827±0.012 0.808±0.039 0.806±0.037 0.802±0.036
BaseSVC / / / /
B-ICB-S 0.805±0.020 0.771±0.031 0.780±0.034 0.771±0.024
IBCB-$\alpha$-1 0.882±0.012 0.870±0.028 0.871±0.033 0.865±0.029
IBCB-$\alpha$-0.6 0.874±0.011 0.863±0.032 0.861±0.038 0.854±0.033
IBCB-$\alpha$-0.8 0.880±0.012 0.868±0.030 0.868±0.034 0.862±0.030
IBCB-$\alpha$-1.2 0.884±0.012 0.872±0.028 0.873±0.032 0.866±0.029
IBCB-$\alpha$-1.5 0.886±0.011 0.872±0.029 0.873±0.033 0.867±0.028
IBCB-$\alpha$-2 0.886±0.012 0.874±0.026 0.875±0.032 0.869±0.028
Algorithm Train Time(sec.) (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB / / / /
B-IRL 36.997±0.803 37.670±1.328 39.910±3.419 37.424±0.563
BaseSVC 116.86±20.63 116.39±26.58 132.18±19.86 117.25±19.47
B-ICB-S 86.365±0.827 87.599±4.489 86.85±1.162 85.713±0.676
IBCB-$\alpha$-1 6.622±2.621 7.992±3.218 7.186±1.577 8.615±4.599
IBCB-$\alpha$-0.6 6.867±2.321 6.802±2.731 7.140±2.044 7.495±4.268
IBCB-$\alpha$-0.8 6.912±2.377 7.913±3.366 7.118±1.762 8.783±5.314
IBCB-$\alpha$-1.2 7.252±2.970 9.823±3.789 7.745±1.758 10.90±6.446
IBCB-$\alpha$-1.5 1.494±0.380 1.453±0.442 1.412±0.359 1.858±0.679
IBCB-$\alpha$-2 1.806±0.465 1.423±0.494 1.620±0.480 1.864±0.686
Table 14: Expert policy parameters’ estimation on synthetic dataset (IBCB’s parameter test). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. Original Expert is SBUCB.
Algorithm Batch Test Log Fitness (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 1 1 1 1
B-IRL 0.828±0.005 0.820±0.013 0.820±0.014 0.818±0.016
BaseSVC 0.791±0.008 0.781±0.020 0.781±0.020 0.777±0.002
B-ICB-S 0.800±0.032 0.767±0.055 0.780±0.045 0.776±0.022
IBCB-$\alpha$-1 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-0.6 0.856±0.011 0.843±0.029 0.842±0.030 0.838±0.027
IBCB-$\alpha$-0.8 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-1.2 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-1.5 0.859±0.011 0.842±0.028 0.843±0.031 0.838±0.027
IBCB-$\alpha$-2 0.858±0.010 0.843±0.027 0.843±0.032 0.839±0.027
Algorithm Batch Test Average Reward (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.641±0.010 0.629±0.028 0.630±0.024 0.622±0.024
B-IRL 0.588±0.015 0.571±0.037 0.572±0.034 0.561±0.034
BaseSVC 0.577±0.017 0.559±0.042 0.559±0.037 0.547±0.037
B-ICB-S 0.581±0.006 0.553±0.024 0.560±0.025 0.552±0.025
IBCB-$\alpha$-1 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-0.6 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-0.8 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-1.2 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-1.5 0.599±0.016 0.580±0.042 0.580±0.040 0.569±0.038
IBCB-$\alpha$-2 0.599±0.016 0.580±0.042 0.580±0.040 0.569±0.038
‘IBCB-$\alpha$-k’ denotes IBCB’s hyper parameter (in Eq.(10)) is k. As shown in Table 13 and Table 14, IBCB’s hyper parameter $\alpha$ will influence several indicators, especially the OL-Fitness and train time, due to the OSQP tolerance level will be relaxed if $\alpha$ is above 1.5. So for IBCB of large $\alpha$ with more relaxed tolerance, $i.e.,$ lower accuracy, training will be faster. Meanwhile, the performance of expert policy’s parameter estimation is similar according to BT-Fitness and BT-AR. By examining the parameters learned by IBCB with different $\alpha$ values, we found that after normalization (by dividing each parameter’s $l_{2}$-norm), the sum of absolute differences in every dimension of the parameters from each pair of two distinct parameters is tightly small. This suggests that one of these parameters can be a result of scaling up or down from another. During BT-data testing, items are recommended based on the parameter under BCB setting without exploration. Therefore, scaling up or down does not influence the recommended item since $\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\hat{\bm{\theta}},\bm{s}_{I}\rangle$ is only determined by the item’s content $\bm{s}_{I}$ when $\hat{\bm{\theta}}$ is fixed with a scaling factor. On the synthetic dataset, the item’s content $\bm{s}_{I}$ is very easy to separate (interval difference of each dimension of $\bm{s}_{I}$ is very obvious through the setting in Appendix B.1), so two distinct parameters with very small sum of absolute differences may share almost the same BT-Fitness and BT-AR.
B.4.4 Results of Ablation Study on ML-100K Dataset
Table 15: Reward & policy parameters’ estimation on ML-100K dataset (using first half of expert’s behavioral evolution history logs). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison.
Algorithm OL-Fitness BT-Fitness BT-AR
SBUCB 0.891±0.006 1 0.193±0.005
B-IRL 0.789±0.017 0.784±0.022 0.189±0.004
BaseSVC / 0.701±0.022 0.186±0.004
B-ICB-S 0.399±0.110 0.325±0.123 0.140±0.025
IBCB (Ours) 0.843±0.019 0.821±0.027 0.191±0.006
Table 16: IBCB’s parameter test on ML-100K dataset. Experts do not need to train since they produce train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm OL-Fitness BT-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 1 0.193±0.005 /
B-IRL 0.830±0.010 0.864±0.008 0.189±0.005 43.669±1.282
BaseSVC / 0.804±0.012 0.189±0.005 162.767±12.135
B-ICB-S 0.746±0.032 0.767±0.028 0.187±0.007 147.474±4.633
IBCB-$\alpha$-1 0.870±0.014 0.884±0.021 0.192±0.006 15.023±3.854
IBCB-$\alpha$-0.6 0.829±0.019 0.884±0.021 0.192±0.006 14.942±3.807
IBCB-$\alpha$-0.8 0.854±0.017 0.884±0.021 0.192±0.006 14.902±3.813
IBCB-$\alpha$-1.2 0.880±0.015 0.880±0.024 0.192±0.006 62.138±59.186
IBCB-$\alpha$-1.5 0.883±0.014 0.876±0.022 0.192±0.006 5.647±0.859
IBCB-$\alpha$-2 0.888±0.010 0.879±0.023 0.192±0.006 6.722±1.641
For ML-100K dataset, Table 15 and Table 16 reports the result of different ablation studies. Results are similar to those on synthetic dataset in most indicators: IBCB can learn from novice expert’s evolution history with better performance compared with other baselines. Note that mean and variance of train time for IBCB-$\alpha$-1.2 is very large. From the OSQP train record, we analysed that this happened due to OSQP failed to convergent for some of the OL-data train logs, so for large $\alpha$, it is recommended to use more relaxed tolerance level in OSQP setting. Note that B-ICB-S’s performance significantly decline when using first half of the train logs in Table 15, this is because when in first half of the real-world train logs, novice expert is still exploring with large steps in a batched manner rather than updating after each step or staying stable with little exploration, which B-ICB-S assumed.
B.4.5 Results of Extra Ablation Study Experiments on ML-100K Dataset
Furthermore, we conducted some extra ablation study experiments on ML-100K to test the real-world task’s performance (e.g. train time, average reward in Batch Test period) improvements against other baselines.
Table 17: Comparison made between original algorithms and baselines using average method when training on ML-100K dataset. Subtable 17(a) is the performance made by original ones, which use whole data to train. Subtable 17(b) is the performance made by algorithms using average method (only baselines, excluding SBUCB and IBCB), which use every episode data training distinct models, and final model uses simple average method to average these models’ parameters.
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 0.193±0.005 /
B-IRL 0.830±0.010 0.189±0.005 43.669±1.282
BaseSVC / 0.189±0.005 162.767±12.135
B-ICB-S 0.746±0.032 0.187±0.007 147.474±4.633
IBCB (Ours) 0.870±0.014 0.192±0.006 15.023±3.854
(a) results of the original experiment
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 0.193±0.005 /
B-IRL 0.835±0.007 0.190±0.005 85.352±9.192
BaseSVC / 0.189±0.004 20.619±0.377
B-ICB-S 0.336±0.076 0.129±0.019 214.288±2.558
IBCB (Ours) 0.870±0.014 0.192±0.006 19.514±4.891
(b) results of averaging $N$ baseline model’s parameters
Table 18: Comparison made between larger batch size $B$ ($B=2000$) and smaller batch size $B$ ($B=500$) when training & testing on ML-100K dataset. Subtable 18(a) is the experiment results obtained when $B=2000$. Subtable 18(b) is the experiment results obtained when $B=500$.
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.896±0.001 0.194±0.005 /
B-IRL 0.816±0.023 0.192±0.005 52.485±1.241
BaseSVC / 0.191±0.004 230.26±12.606
B-ICB-S 0.685±0.028 0.188±0.006 144.697±7.972
IBCB (Ours) 0.851±0.016 0.196±0.007 15.083±2.036
(a) results of setting batch size $B=2000$
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.89±0.012 0.193±0.006 /
B-IRL 0.845±0.007 0.190±0.006 65.646±1.378
BaseSVC / 0.189±0.005 266.538±12.153
B-ICB-S 0.716±0.039 0.187±0.007 138.616±8.982
IBCB (Ours) 0.862±0.019 0.191±0.005 26.255±10.370
(b) results of setting batch size $B=500$
Firstly, since we assume that training period can be described as a period containing $N$ episodes’ training data, we want to know how this $N$’s setting influence the performance of baselines. Specifically, we divided the total $N$ episodes to individually train corresponding IRL or BC models. Finally, we averaged the parameters of the $N$ models obtained from each episode’s training to form the final test model. In this setup, the training time equates to the cumulative time for training $N$ models (disregarding multi-threading, assuming that each time we collect data from an episode, we train a model sequentially). The experiment was conducted using the ML-100K dataset, as we aimed to investigate the practical feasibility. Detailed experimental results can be found in Table 17. From the experimental results, it is evident that training IRL or BC models separately for each episode does yield certain differences in comparison to the unified training approach. The B-IRL model demonstrates a slight improvement in performance (e.g. OL-Fitness), albeit with a significant increase in training time. This could be attributed to the necessity of restarting training each time due to the inability to iteratively utilize previously acquired parameters when new data is collected. Of particular interest is the significant decline in performance for B-ICB. This discrepancy is likely linked to the inconsistency in retaining the original expert model’s (i.e. SBUCB) parameters within each episode, which contradicts B-ICB’s assumption of evolving data within every single step. Additionally, due to the necessity for repeated restarts during training, the overall training time becomes extended. As for BaseSVC, the reduced training sample size in each episode leads to notably shorter training time for individual runs (attributed to its near-quadratic time complexity). While this reduces overall training time, the efficacy is comparatively diminished compared to the direct training using the entire dataset. Ultimately, even in comparison with the models subjected to parameter averaging, the performance of IBCB remains superior overall. This underscores the effectiveness of using episodes as evolutionary segments in IBCB’s training process in real-world datasets.
Secondly, we wanted to work out the influence of batch size $B$’s setting. Throughout this ablation study experiment, we controlled the total number of training and testing samples and fixed the context, ensuring that the product of the total number of episodes $N$ and the batch size $B$ within each episode remains constant. Based on this foundation, we adjusted the batch size $B$. Specifically, the original values were $N=20$ and $B=1000$. Extra experiments included scenarios with larger $B$, denoted as $N=10$ and $B=2000$ , with corresponding results displayed in Table 18(a). Likewise, there were scenarios with smaller $B$, represented as $N=40$ and $B=500$, with results exhibited in Table 18(b). From the results of these additional experiments and comparisons with the original experiments, we deduced that, while controlling the total number of training and testing samples and context, a larger $B$ yields shorter overall training time and better average reward performance on the test set. However, the fitting to the environment is reduced due to the decreased value of $N$, implying that more data for training comprises novice expert’s evolutionary history, hence fostering more exploration. With reduced $N$ and increased $B$, the frequency of evolution decreases, making adaptation to the original training data easier, thereby shortening training time. In such cases, the estimations of various baselines and IBCB regarding the environment are not as precise as those under smaller $B$. Appropriately crafting exploration items can result in higher rewards. Conversely, a smaller $B$ leads to a better fit with the environment, albeit resulting in a longer training time and inferior average reward performance on the test set, which aligns with the experimental results.